The newly installed elevator of a building weighs 4000 lbs. During the test run, it rises with an acceleration of 4 ft/sec2 . Find the tension in the supporting cable of the elevator


1
Expert's answer
2021-03-31T09:38:13-0400

Let's apply the Newton's Second Law of Motion:


"T-mg=ma,""T=m(g+a)=4000\ lbs\cdot(32.17\ \dfrac{ft}{s^2}+4\ \dfrac{ft}{s^2})=144680\ \dfrac{ft\cdot lbs}{s2}."